Sketch the graph of the resulting function. 2. Solve a +x = 8(t - ) - 8(t-2r), z(0) = 0, z(0) = 1. Sketch the graph of the
Answer #1

Que 2. altx aslt-a) a s (t-20) 1 4020 We will solving using taplace transfrom &(o)-1 2{x} = 82x8) - sxo) - 260) L {s(t-a) fQue 3 (a) y -44 +44.0 Characteristics eq. m2-4m +400 (m ) 2=0 m =2,2 ylt) = c est + Ctet (6) £zy -4ty +4y=0 -0 It is of tX=AX Aus-a) A= of above matrix, Find eigen value & eigen vector det (A-AI) =0 -4 3 (Ind) (37)=0 d=1,3 for da 1 CA-I) X=0 -Ausb) A = 5 det (A-d I) 20 CI-d) (-3d) +500 -3rd+30 +2 +50 72 +24+2 -0. da - 1 ti da-i ti for CA-AI) to ri-(-1+i) 1 - 5 -3-(-